116% 0f t is 50
150% of 90
what percent of 40 is 25
find 500% of 12
what percent of 25 is 17
85% of z is 106,250.what is z

Answer by rfer(16322)

50*1.16=58
25/40=0.625=62.5%
0.85z=106250
z=106250/0.85
z=125,000
you try the others